OpenStudy (anonymous):

Cashews cost $5 per pound and peanuts cost $1.50 per pound. How many pounds of cashews should be mixed with peanuts to get a mixture of 32 pounds that will cost $4 per pound? Round to the nearest tenth.

OpenStudy (anonymous):

p + c = 32 5c + 1.5p = 4 johnweldon....is that the correct set up ?

OpenStudy (johnweldon1993):

hmmm I thought that same thing...but I think the final cost would be 5c + 1.5p = (32 * 4) since it states that the final mixture would be 32 pounds that cost $4 a pound... just my thought...what do you think...?

OpenStudy (anonymous):

oh...you might be right....this reminds me of the question I had last night...lol

OpenStudy (johnweldon1993):

I know I saw it and was like...oh no not another one of these! lol

OpenStudy (anonymous):

just some " nutty " questions

OpenStudy (johnweldon1993):

-_- good one lol :P

OpenStudy (anonymous):

5c + 1.5p = 32 * 4 5c + 1.5p = 128 where do I go from here....division ?

OpenStudy (johnweldon1993):

Well no...now you solve the system... c + p = 32 5c + 1.5p = 128

OpenStudy (anonymous):

oh.....thats easy....I am trying to make things too hard

OpenStudy (anonymous):

corrinee can probably do this

OpenStudy (johnweldon1993):

yeah don't do that! lol then we get *your question from last night* lol can you solve this system @corrinee ?

OpenStudy (anonymous):

I am gonna take off now...I am hungry....pizza sound good...see you peeps later

OpenStudy (johnweldon1993):

haha so long @kelliegirl33 :)